Required information [The following information applies to the questions displayed below.] Cane Company manufactures two prod14. Assume that Canes customers would buy a maximum of 96,000 units of Alpha and 76,000 units of Beta. Also assume that the15. Assume that Canes customers would buy a maximum of 96,000 units of Alpha and 76,000 units of Beta. Also assume that the
Answer #1

Contribution margin per pound

Alpha Beta
Contribution margin per unit 87 66
Pound per unit 6 3
Contribution margin per pound 14.5 22

13) Unit produced

Alpha Beta
Unit produced (246000-228000)/6 = 3000 76000

14) Total Contribution margin = (3000*87+76000*66) = $5277000

15) Highest price = 7+14.50 = 21.50